# Spin Chern number in altermagnets

### v1

@@ Introduction
Altermagnetism has emerged as a novel class of collinear magnetic phase distinguished by a unique breaking of time-reversal symmetry ($\mathbb{T}$) in reciprocal space, despite of having a compensated magnetic order in real space. Unlike conventional collinear antiferromagnets (AFMs), altermagnets exhibit a distinct electronic band structure due to their sublattices of opposite spin being connected not through simple translation or inversion, but through non-trivial rotations 1, 2, 3. For example, the combination of four-fold rotation with time-reversal symmetry ($C^{z}_{4}\mathbb{T}$) was identified in the metallic RuO2, this one being the first $d$-wave altermagnetic material discovered 4. In this system, anisotropic spin charge densities lead to considerable spin-splitting in the electronic bands, which is independent of relativistic spin-orbit coupling (SOC) interactions. This spin-splitting produce unconventional magnetic responses, such as the anomalous Hall 4, 5, 6 , spin-filter 7, 8 and the magneto-optical Kerr effect 9, which make altermagnets an interesting topic for both theoretical  10, 11, 12, 13, 14, 15, 16, 17 and experimental  18, 19, 20, 21, 22 investigation in condensed matter physics.
@@
On the other hand, the topological materials can be classified into categories such as topological insulators and topological semimetals 23, 24, 25. The latter can defined by the presence of symmetry protected Weyl points (WPs) close to the Fermi level. These WPs usually appear when spin degeneracy is lifted through mechanisms like strong SOC in inversion or time-reversal symmetry breaking systems 26, 27. While traditional AFMs cannot support WPs due to their spin-degenerate bands, altermagnets are expected to exhibit WPs naturally due to their spin-split bands 1, 2. These unique properties have generated interest in the combination of altermagnetic and topological properties which are protected by symmetry operations.
@@
However, defining topological invariants in insulating or semimetallic altermagnets poses a challenge, particularly due to the absence of time-reversal symmetry, which renders the $\mathbb{Z}_{2}$ index undefined 28, 29. Additionally, the $C^{z}_{4}\mathbb{T}$ symmetry causes the Chern number to vanish in the entire Brillouin zone (BZ), which also restricts its potential as a possible topological indicator 30, 31. Recently, the $C^{z}_{4}\mathbb{T}$ symmetry has been employed to divide the BZ into two regions, allowing for the calculation of a half-quantized Chern number with opposite signs 32.
@@
In this work, we have employed equivariant K-theory of magnetic groups to establish a robust $C^{z}_{4}\mathbb{T}$ topological invariant for characterizing altermagnetic systems. Using K-theory and Hamiltonian models, we demonstrates that the spin Chern number (SCN) 33 can be a topological index for classifying 2D and 3D insulating and semimetallic altermagnets with $C^{z}_{4}\mathbb{T}$ symmetry-protected states. We prove that the SCN corresponds to the half-quantized Chern number assigned to the divided BZ, and therefore the SCN becomes a clear indicator to identify topological altermagnetic phases. Finally, the topologically protected 2D insulator and 3D Weyl semimetal altermagnets are predicted through spin topology characterization 34, 35.
@@
Figure 1: (a) Spin splitting of the electronic band structure of the two-dimensional (2D) altermagnet topological insulator (TI) derived from the Hamiltonian in Eqn. (1), with $G$=1.05. (b) Spin projection at the top of the valence band and (c) Berry curvature for the 2D altermagnet TI with a mass term $M_{0}=0.4$. (d) Phase diagram of spin-up Chern number $c^{\uparrow}$ and the integral of Berry curvature $\Omega_{1}$ for the region $R_{1}^{\uparrow}$ as a function of the mass term $M_{0}$. Non-trivial and trivial phases are classified with the variation of the Chern numbers. (e) Spin projection at the top of the valence band ($M_{0}=1.7$), illustrating the nodal lines at $k_{x}=\pm k_{y}$ where a change in spin value occurs. (f) Berry curvature and (g) Spin Berry curvature for the case of $M_{0}=1.7$. All the Hamiltonian parameters are in $eV$. The integration regions $R_{1}^{\uparrow}$ and $R_{1}^{\downarrow}$ are indicated in panel (f). The Fermi energy is set to zero.
We consider a four band model on a square lattice in which unit-cell contains the following orbitals: $d_{xy\uparrow}$, $\frac{1}{\sqrt{2}}(d_{xz\uparrow}-id_{yz\uparrow})$, $d_{xy\downarrow}$ and $\frac{1}{\sqrt{2}}(d_{xz\downarrow}-id_{yz\downarrow})$. In this model, a 2D altermagnetic Hamiltonian similar to the Bernevig-Hughes-Zhang (BHZ) 36 in momentum space, has been proposed by  Ma and Jia 37, taking the following form:
@@
Here, $M_{1}$ and $M_{2}$ represent the mass terms for each spin channel, while $M_{0}$ denotes the on-site atomic potential which can vary with the value of the local magnetic moment. The parameters $A_{0}$ and $B_{0}$ describe the hopping amplitudes between $d$ orbitals and we will set these to $1$ eV for convenience. The constant $G$ introduces a symmetry breaking between the $x$ and $y$ directions within each spin block, making the Hamiltonian characteristic of a $d$-wave altermagnetic 2D system 37. We set the constant $G$ to be a number slightly greater than $1$ (in $eV$):
$G>1\ \ \&\ \ G\sim 1.$ (6)
@@
Due to the spin character changing inside each of the valence bands, as shown in Fig. 1e, this spin texture induce nodal lines in the altermagnetic systems, as predicted by several works 38, 39, 40, 41. The valence bands intersect along the nodal lines at $k_{x}=\pm k_{y}$ in reciprocal space. When no additional intersections occur, the upper and lower valence bands split into two connected regions, as depicted in Fig. 1e. This is the case when $|M_{0}|>G^{2}$.
@@
Therefore the integral of the Berry curvature for each spin region also permits to determine the spin Chern number. To calculate the Chern and spin Chern numbers, we have constructed the Hamiltonian model using the pythtb code 42.
@@
For values of $M_{0}$ bigger than $G^{2}$, the valence bands separate, except on the lines $k_{x}=\pm k_{y}$, and the value of the integral of the Berry curvature is $\tfrac{1}{2}$. This result is consistent with the half-quantized Berry curvature predicted for two-dimensional $C_{4z}\mathbb{T}$-symmetric topological semimetals 32.
@@
The topological character of this 2D altermagnet can be distinguished by measuring the quantized spin Hall response, which is determined by the spin Chern number. As we have demonstrated, the spin Chern number can be a robust topological invariant for characterizing 2D altermagnetic phases with $C_{4z}\mathbb{T}$ symmetry, which allows a clean separation of the spin channels for each valence bands. This invariant can be used to predict the topological phases of the $d$-wave 2D altermagnetic materials, such as FeSe43, as well as those recently proposed by  Bai et al. 44
@@
In Figure 2, we present the results of the 3D Hamiltonian from Eqn. (20) under conditions of negligible SOC, with $D_{0}$ and $C_{0}$ set to zero. Figure 2a) shows one of the Weyl points (WP) in the 2D energy spectrum, specifically located within an particular $k_{z}$ plane. The positions of the all WPs in the reciprocal space are shown in Figure 2b), indicating that only two WPs are found along the $k_{z}$ axis
@@
To validate our results, we compute the surface states using the iterative Green’s function method in a plane parallel to the $z$-axis 45. Figure 2e) shows a Fermi arc connecting the WPs in the surface energy spectrum of the $xz$ plane, providing further evidence to detect this 3D altermagnetic topological semimetal phase. Recently, Fermi arcs have been both theoretically predicted and experimentally observed in the room-temperature 3D altermagnet candidate CrSb 46.
Now, we will include the SOC into the 3D Hamiltonian described by Eqn. (20). In this case, whenever the term $D_{0}$ is zero, the Hamiltonian preserves both $C_{4}\mathbb{T}$ and $I$. When $D_{0}$ is not zero, inversion is broken and $C_{4}\mathbb{T}$ is preserved. Here the valence bands are non-degenerate and therefore $I\mathbb{T}$ is not preserved. The energy bands are separated by either making $G\neq 1$ or $C+iD\neq 0$.
@@
The spin spectrum becomes non-degenerate except on spin Weyl points localized along the axis $(k_{x},k_{y})=(0,0)$ for $G<M_{0}<2+G$, $(k_{x},k_{y})=(0,\pi)$ for $(k_{x},k_{y})=(\pi,0)$ $-2+G<M_{0}<2-G$ and $(k_{x},k_{y})=(\pi,\pi)$ for $-2-G<M_{0}<-G$. The position in $k_{z}$ of the spin Weyl points (SWPs) solve the equations $\cos(k_{z})=M_{0}-1-G$, $\cos(k_{z})=M_{0}\pm 1\mp G$ and $\cos(k_{z})=M_{0}+1+G$ respectively.
In Figure 3a) it is shown the locations of WPs and SWPs 34 for the parameters $G=1.05$ and $M_{0}=1.8$ in the 3D Hamiltonian described by Eqn. (20). The SOC parameters in the Hamiltonian are set to $C_{0}$=$1.0$ and $D_{0}$=$0.3$, with all Hamiltonian parameters in $eV$. Notably, a total of eight WPs and two SWPs are identified in reciprocal space, with the SWPs located along the $k_{z}$ axis, as predicted. From Figure 3b), it is noted that the spin Chern number changes when the $k_{z}$ planes cross any of the WPs and SWPs. In this case, the $C$ term splits the $k_{z}$ location of the energy and spin WPs, as is reflected by the spin-up Chern number ($c^{\uparrow}$) calculation as a function of the $k_{z}$ plane variable.
@@ III Topological invariants of $C_{4}^{z}\mathbb{T}$
In this section we will determine the bulk invariants in the of the 2D and 3D torus with the action of $C_{4}^{z}\mathbb{T}$. For this we need some results on equivariant K-theory of magnetic groups that will appear in forthcoming publications 47, 48.
@@ III.1 Bulk invariants of the magnetic K-theory
and will be called the magnetic $G$-equivariant K-theory of $X$. The complete K-theory groups $\mathcal{K}^{*}_{G}(X)$ were firstly proposed by Karoubi 49, further developed and related to electronic properties of crystals by Freed and Moore 50, and they have been extensively studied by other authors 51, 47, 48. The definition of the higher magnetic equivariant K-theory groups is the standard one, namely
@@
We will make an abuse of notation and we will not differentiate in the notation the case on which spin orbit coupling (SOC) is present no the one on which is absent. Whenever SOC is present, the appropriate magnetic K-theory groups need to be twisted 51. Alternatively, we will assume that the magnetic group is a $\mathbb{Z}_{2}$ central extension of the magnetic point group. We will simply assume that in the presence of SOC the time reversal operator $\mathbb{T}$ squares to $-1$ and all rotations after a whole turn multiply by $-1$. The $G$ action on the fibers of the complex vector bundle $E$ need to take this information into account.
@@
The three copies of $\mathbb{Z}_{2}$ on the 3D torus are called the weak invariants, while the fourth copy is the strong invariant for topological insulators 52.
@@
Since $G_{0}$ acts complex linearly, the restricted K-theory groups are the standard (twisted) equivariant complex K-theory groups53. These will be denoted with the symbol $K$ for unitary K-theory. The curly $\mathcal{K}$ will be kept for magnetic K-theory.
When applied to a point, this restriction map takes an irreducible magnetic representation of the group $G$ (or corepresentation in Wigner’s notation 54) and returns the complex irreducible representations that make the magnetic one:
@@
Here $\widehat{U}$ denotes the conjugate representation of $U$ and it is defined as the complex conjugate representation of the conjugate $G_{0}$ action; see Eqn. (36). The important things to note are the following 54:
@@
recall that $aga^{-1}$ is also in $G_{0}$ since $G_{0}$ is normal in $G$. The bundle $\widehat{F}$ satisfies several properties 47:
@@
  • When $X$ is a point, $\widehat{V}$ denotes the conjugate representation of $V$ defined by Wigner 54.
We claim the following results47:
@@
III.2 $\mathbb{Z}_{2}$ bulk invariant for $C_{4}\mathbb{T}$ on 2D torus
Denote by $\mathcal{K}^{*}_{C_{4}\mathbb{T}}(T^{2})$ the K-theory groups of the 2D torus with the action of the group generated by $C_{4}\mathbb{T}$. Here we are assuming that we have SOC and therefore $(C_{4}\mathbb{T})^{4}=-1$. We are interested in finding the bulk invariants for this magnetic group and we will make use of the isomorphism of Eqn. (40).
Let us start understanding the magnetic K-theory groups over a point, namely the magnetic representations. There is only one non-trivial isomorphism class of irreducible representations of the group $\langle C_{4}\mathbb{T}\rangle$ and it is parametrized by the following $2\times 2$ complex matrices:
$C_{4}\mathbb{T}\mapsto\left(\begin{matrix}0&i\\ -1&0\end{matrix}\right)\mathbb{K},\ \ \ (C_{4}\mathbb{T})^{2}\mapsto\left(\begin{matrix}-i&0\\ 0&i\end{matrix}\right),$ (41)
where $\mathbb{K}$ denotes complex conjugation; denote this representation $V$. This representation is of complex type (since the restriction to $C_{2}$ splits into two non-isomorphic irreducible representations), and therefore the K-theory groups of a point are:
$\mathcal{K}^{even}_{C_{4}\mathbb{T}}(*)\cong\mathbb{Z},\ \ \ \mathcal{K}^{odd}_{C_{4}\mathbb{T}}(*)\cong 0,$ (42)
where $\mathbb{Z}$ parametrizes the number of copies of the only irreducible representation $V$ of $\langle C_{4}\mathbb{T}\rangle$.
For the group $C_{2}$ we have the following:
$\mathcal{K}^{even}_{C_{2}}(*)\cong\mathbb{Z}\oplus\mathbb{Z},\ \ \ \mathcal{K}^{odd}_{C_{2}}(*)\cong 0.$ (43)
where $\mathbb{Z}\oplus\mathbb{Z}$ parametrizes the number of irreducible representations $L$ and $\widehat{L}$ of $C_{2}$. These representations are 1-dimensional and the generator of $C_{2}$ acts on $L$ by multiplication by $i$ and on $\widehat{L}$ by $-i$. Here we need to recall that $(C_{2})^{2}=-1$.
Since the only magnetic representation of $C_{4}\mathbb{T}$ is of complex type, the restriction map defines an isomorphism at the level of representations:
$\mathcal{K}^{0}_{C_{4}\mathbb{T}}(*)$ $\stackrel{{\scriptstyle\cong}}{{\to}}{K}^{0}_{C_{2}}(*)^{\mathbb{Z}_{2}}$ (44)

### v2

@@ Introduction
Altermagnetism 1, 2, 3, 4, 5, 6 has emerged as a novel class of collinear magnetic phase distinguished by a unique breaking of time-reversal symmetry ($\mathbb{T}$) in reciprocal space, despite of having a compensated magnetic order in real space. Unlike conventional collinear antiferromagnets (AFMs), altermagnets exhibit a distinct electronic band structure due to their sublattices of opposite spin being connected not through simple translation or inversion, but through non-trivial rotations 7, 8, 9, 10. For example, the combination of four-fold rotation with time-reversal symmetry ($C^{z}_{4}\mathbb{T}$) was identified in the metallic RuO2, this one being the first $d$-wave altermagnetic material discovered 1, 11. In this system, anisotropic spin charge densities lead to considerable spin-splitting in the electronic bands, which is independent of relativistic spin-orbit coupling (SOC) interactions. This spin-splitting produce unconventional magnetic responses, such as the anomalous Hall 12, 13, 14, spin-filter 15, 16 and the magneto-optical Kerr effect 17, which make altermagnets an interesting topic for both theoretical  18, 19, 20, 21, 22, 23, 24, 25, 26 and experimental  27, 28, 29, 30, 31 investigation in condensed matter physics.
@@
On the other hand, the topological materials can be classified into categories such as topological insulators and topological semimetals 32, 33, 34. The latter can defined by the presence of symmetry protected Weyl points (WPs) at the Fermi level. These WPs usually appear when spin degeneracy is lifted through mechanisms like strong SOC in inversion or time-reversal symmetry breaking systems 35, 36. While traditional AFMs cannot support WPs due to their spin-degenerate bands, altermagnets are expected to exhibit WPs naturally due to their spin-split bands 2, 3. These unique properties have generated interest in the combination of altermagnetic and topological properties which are protected by symmetry operations.
@@
However, defining topological invariants in insulating or semimetallic altermagnets poses a challenge, particularly due to the absence of time-reversal symmetry, which renders the $\mathbb{Z}_{2}$ index undefined 37, 38. Additionally, the $C^{z}_{4}\mathbb{T}$ symmetry causes the Chern number to vanish in the entire Brillouin zone (BZ), which also restricts its potential as a possible topological indicator 39, 40. Recently, the $C^{z}_{4}\mathbb{T}$ symmetry has been employed to divide the BZ into two regions, allowing for the calculation of a half-quantized Chern number with opposite signs 41.
@@
In this work, we have employed equivariant K-theory of magnetic groups to establish a robust $C^{z}_{4}\mathbb{T}$ topological invariant for characterizing altermagnetic systems. Using K-theory and Hamiltonian models, we demonstrates that the spin Chern number (SCN) 42 can be a topological index for classifying 2D and 3D insulating and semimetallic altermagnets with $C^{z}_{4}\mathbb{T}$ symmetry-protected states. We prove that the SCN corresponds to the half-quantized Chern number assigned to the divided BZ, and therefore the SCN becomes a clear indicator to identify topological altermagnetic phases. Finally, the topologically protected 2D insulator and 3D Weyl semimetal altermagnets are predicted through spin topology characterization 43, 44, 45.
@@
Figure 1: (a) Spin splitting of the electronic band structure of the two-dimensional (2D) altermagnet topological insulator (TI) derived from the Hamiltonian in Eqn. (1), with $G$=1.05. (b) Spin projection at the top of the valence band and (c) Berry curvature for the 2D altermagnet TI with a mass term $M_{0}=0.4$. (d) Phase diagram of spin-up Chern number $c^{\uparrow}$ and the integral of Berry curvature $\Omega_{1}$ over half of the Brillouin zone $R$, as shown in the panel, as a function of the mass term $M_{0}$. Non-trivial and trivial phases are classified with the variation of the Chern numbers. Note that the value of the Chern number depends on the radius of the nodal line for $M_{0}<G^{2}$. (e) Spin projection at the top of the valence band ($M_{0}=1.7$), illustrating the nodal lines at $k_{x}=\pm k_{y}$ where a change in spin value occurs. (f) Berry curvature and (g) Spin Berry curvature for the case of $M_{0}=1.7$. All the Hamiltonian parameters are in $eV$. The integration regions $R_{1}^{\uparrow}$ and $R_{1}^{\downarrow}$, used in the calculations of $c_{1}^{\uparrow}$ and $c_{1}^{\uparrow}$ in Eqn. 17, are shown in panels (c) and (f). The Fermi energy is set to zero.
We consider a four band model on a square lattice in which unit-cell contains the following orbitals: $d_{xy\uparrow}$, $\frac{1}{\sqrt{2}}(d_{xz\uparrow}-id_{yz\uparrow})$, $d_{xy\downarrow}$ and $\frac{1}{\sqrt{2}}(d_{xz\downarrow}-id_{yz\downarrow})$. In this model, a 2D altermagnetic Hamiltonian similar to the Bernevig-Hughes-Zhang (BHZ) 46 in momentum space, has been proposed by Ma and Jia47, taking the following form:
@@
Here, $M_{1}$ and $M_{2}$ represent the mass terms for each spin channel, while $M_{0}$ denotes the on-site atomic potential which can vary with the value of the local magnetic moment. The parameters $A_{0}$ and $B_{0}$ describe the hopping amplitudes between $d$ orbitals and we will set these to $1$ eV for convenience. The constant $G$ introduces a symmetry breaking between the $x$ and $y$ directions within each spin block, making the Hamiltonian characteristic of a $d$-wave altermagnetic 2D system when G $\neq$ 1 47. We set the constant $G$ to be a number slightly shifted from $1$ (in $eV$):
$G\sim 1.$ (6)
@@
Due to the spin character changing inside each of the valence bands, as shown in Fig. 1e, this spin texture induce nodal lines in the altermagnetic systems, as predicted by several works 48, 49, 50, 51. The valence bands intersect along the nodal lines at $k_{x}=\pm k_{y}$ in reciprocal space. When no additional intersections occur, the upper and lower valence bands split into two connected regions, as depicted in Fig. 1e. This is the case when $|M_{0}|>G^{2}$.
@@
Therefore the integral of the Berry curvature for each spin region also permits to determine the spin Chern number. To calculate the Chern and spin Chern numbers, we have constructed the Hamiltonian model using the pythtb code 52.
@@
For values of $M_{0}$ bigger than $G^{2}$, the valence bands separate, except on the lines $k_{x}=\pm k_{y}$, and the value of the integral of the Berry curvature is $\tfrac{1}{2}$. This result is consistent with the half-quantized Berry curvature predicted for two-dimensional $C_{4}^{z}\mathbb{T}$-symmetric topological semimetals 41.
@@
The topological character of this 2D altermagnet can be distinguished by measuring the quantized spin Hall response, which is determined by the spin Chern number. As we have demonstrated, the spin Chern number can be a robust topological invariant for characterizing 2D altermagnetic phases with $C_{4}^{z}\mathbb{T}$ symmetry, which allows a clean separation of the spin channels for each valence bands. This invariant can be used to predict the topological phases of the $d$-wave 2D altermagnetic materials, such as FeSe53, as well as those recently proposed by Bai et al. 9
@@
In Figure 2, we present the results of the 3D Hamiltonian from Eqn. (20) under conditions of negligible SOC, with $D_{0}$ and $C_{0}$ set to zero. Figure 2a) shows one of the Weyl points (WP) in the 2D energy spectrum, specifically located within an particular $k_{z}$ plane. The positions of the all WPs in the reciprocal space are shown in Figure 2b), indicating that only two WPs are found along the $k_{z}$ axis. In agreement with our findings, a recent study 51 predicts Weyl nodal lines in a 3D altermagnetic system obtained by extending a 2D altermagnetic model.
@@
To validate our results, we compute the surface states using the iterative Green’s function method in a plane parallel to the $z$-axis 54. Figure 2e) shows a Fermi arc connecting the WPs in the surface energy spectrum of the $xz$ plane, providing further evidence to detect this 3D altermagnetic topological semimetal phase. Recently, Fermi arcs have been both theoretically predicted and experimentally observed in the room-temperature 3D altermagnet candidate CrSb 55.
Now, we will include the SOC into the 3D Hamiltonian described by Eqn. (20). In this case, whenever the term $D_{0}$ is zero, the Hamiltonian preserves both $C_{4}^{z}\mathbb{T}$ and $I$. When $D_{0}$ is not zero, inversion is broken and $C_{4}^{z}\mathbb{T}$ is preserved. Here the valence bands are non-degenerate and therefore $I\mathbb{T}$ is not preserved. The energy bands are separated by either making $G\neq 1$ or $C+iD\neq 0$.
@@
The spin spectrum becomes non-degenerate except on spin Weyl points43 (SWPs) localized along the axis $(k_{x},k_{y})=(0,0)$ for $G<M_{0}<2+G$, $(k_{x},k_{y})=(0,\pi)$ for $(k_{x},k_{y})=(\pi,0)$ $-2+G<M_{0}<2-G$ and $(k_{x},k_{y})=(\pi,\pi)$ for $-2-G<M_{0}<-G$. The position in $k_{z}$ of the spin Weyl points solve the equations $\cos(k_{z})=M_{0}-1-G$, $\cos(k_{z})=M_{0}\pm 1\mp G$ and $\cos(k_{z})=M_{0}+1+G$ respectively.
In Figure 3a) it is shown the locations of WPs and SWPs 43 for the parameters $G=1.05$ and $M_{0}=1.8$ in the 3D Hamiltonian described by Eqn. (20). The SOC parameters in the Hamiltonian are set to $C_{0}$=$1.0$ and $D_{0}$=$0.3$, with all Hamiltonian parameters in $eV$. Notably, a total of eight WPs and two SWPs are identified in reciprocal space, with the SWPs located along the $k_{z}$ axis, as predicted. From Figure 3b), it is noted that the spin Chern number changes when the $k_{z}$ planes cross any of the WPs and SWPs. In this case, the $C$ term splits the $k_{z}$ location of the energy and spin WPs, as is reflected by the spin-up Chern number ($c^{\uparrow}$) calculation as a function of the $k_{z}$ plane variable.
@@ III Topological invariants of $C_{4}^{z}\mathbb{T}$
In this section we will determine the bulk invariants in the of the 2D and 3D torus with the action of $C_{4}^{z}\mathbb{T}$. For this we need some results on equivariant K-theory of magnetic groups that will appear in forthcoming publications 56, 57.
@@ III.1 Bulk invariants of the magnetic K-theory
and will be called the magnetic $G$-equivariant K-theory of $X$. The complete K-theory groups $\mathcal{K}^{*}_{G}(X)$ were firstly proposed by Karoubi 58, further developed and related to electronic properties of crystals by Freed and Moore 59, and they have been extensively studied by other authors 60, 56, 57. The definition of the higher magnetic equivariant K-theory groups is the standard one, namely
@@
We will make an abuse of notation and we will not differentiate in the notation the case on which spin orbit coupling (SOC) is present no the one on which is absent. Whenever SOC is present, the appropriate magnetic K-theory groups need to be twisted 60. Alternatively, we will assume that the magnetic group is a $\mathbb{Z}_{2}$ central extension of the magnetic point group. We will simply assume that in the presence of SOC the time reversal operator $\mathbb{T}$ squares to $-1$ and all rotations after a whole turn multiply by $-1$. The $G$ action on the fibers of the complex vector bundle $E$ need to take this information into account.
@@
The three copies of $\mathbb{Z}_{2}$ on the 3D torus are called the weak invariants, while the fourth copy is the strong invariant for topological insulators 38.
@@
Since $G_{0}$ acts complex linearly, the restricted K-theory groups are the standard (twisted) equivariant complex K-theory groups61. These will be denoted with the symbol $K$ for unitary K-theory. The curly $\mathcal{K}$ will be kept for magnetic K-theory.
When applied to a point, this restriction map takes an irreducible magnetic representation of the group $G$ (or corepresentation in Wigner’s notation 62) and returns the complex irreducible representations that make the magnetic one:
@@
Here $\widehat{U}$ denotes the conjugate representation of $U$ and it is defined as the complex conjugate representation of the conjugate $G_{0}$ action; see Eqn. (36). The important things to note are the following 62:
@@
recall that $aga^{-1}$ is also in $G_{0}$ since $G_{0}$ is normal in $G$. The bundle $\widehat{F}$ satisfies several properties 56:
@@
  • When $X$ is a point, $\widehat{V}$ denotes the conjugate representation of $V$ defined by Wigner 62.
We claim the following results56:
@@
III.2 $\mathbb{Z}_{2}$ bulk invariant for $C_{4}^{z}\mathbb{T}$ on 2D torus
Denote by $\mathcal{K}^{*}_{C_{4}^{z}\mathbb{T}}(T^{2})$ the K-theory groups of the 2D torus with the action of the group generated by $C_{4}^{z}\mathbb{T}$. Here we are assuming that we have SOC and therefore $(C_{4}^{z}\mathbb{T})^{4}=-1$. We are interested in finding the bulk invariants for this magnetic group and we will make use of the isomorphism of Eqn. (40).
Let us start understanding the magnetic K-theory groups over a point, namely the magnetic representations. There is only one non-trivial isomorphism class of irreducible representations of the group $\langle C_{4}^{z}\mathbb{T}\rangle$ and it is parametrized by the following $2\times 2$ complex matrices:
$C_{4}^{z}\mathbb{T}\mapsto\left(\begin{matrix}0&i\\ -1&0\end{matrix}\right)\mathbb{K},\ \ \ (C_{4}^{z}\mathbb{T})^{2}\mapsto\left(\begin{matrix}-i&0\\ 0&i\end{matrix}\right),$ (41)
where $\mathbb{K}$ denotes complex conjugation; denote this representation $V$. This representation is of complex type (since the restriction to $C_{2}^{z}$ splits into two non-isomorphic irreducible representations), and therefore the K-theory groups of a point are:
$\mathcal{K}^{even}_{C_{4}^{z}\mathbb{T}}(*)\cong\mathbb{Z},\ \ \ \mathcal{K}^{odd}_{C_{4}^{z}\mathbb{T}}(*)\cong 0,$ (42)
where $\mathbb{Z}$ parametrizes the number of copies of the only irreducible representation $V$ of $\langle C_{4}^{z}\mathbb{T}\rangle$.
For the group $C_{2}^{z}$ we have the following:
$\mathcal{K}^{even}_{C_{2}^{z}}(*)\cong\mathbb{Z}\oplus\mathbb{Z},\ \ \ \mathcal{K}^{odd}_{C_{2}^{z}}(*)\cong 0.$ (43)
where $\mathbb{Z}\oplus\mathbb{Z}$ parametrizes the number of irreducible representations $L$ and $\widehat{L}$ of $C_{2}^{z}$. These representations are 1-dimensional and the generator of $C_{2}^{z}$ acts on $L$ by multiplication by $i$ and on $\widehat{L}$ by $-i$. Here we need to recall that $(C_{2}^{z})^{2}=-1$.
Since the only magnetic representation of $C_{4}^{z}\mathbb{T}$ is of complex type, the restriction map defines an isomorphism at the level of representations:
$\mathcal{K}^{0}_{C_{4}^{z}\mathbb{T}}(*)$ $\stackrel{{\scriptstyle\cong}}{{\to}}{K}^{0}_{C_{2}^{z}}(*)^{\mathbb{Z}_{2}}$ (44)